# Token Pruning in Audio Transformers: Optimizing Performance and Decoding Patch Importance

### v1

@@ III-C Correlation Between Attention and Statistical Features
III-D Impact of removing tokens from low/high-intensity regions
We assess the impact of discarding tokens belonging to low- (C1, C2: ) or high-intensity (C4, C5: ) groups on accuracy during inference. In this test, the tokens belonging to these groups are pruned after being processed by specific blocks. In Table IV, and indicate the block index at which low- and high-intensity token groups are removed, respectively. High-intensity tokens contribute significantly to model performance. In addition, low-intensity tokens become more important for general audio classification tasks than for speech tasks. AudioMAE shows greater robustness to the loss of both token types across all datasets compared to AST.
III-E Inspection of retained patches’ statistics
We visualize the relationship between the and of signals in patches for each dataset in Fig. 6. The vertical lines on low-mean patches in the input Mel-spectrograms of ESC-50 and SPC-2 indicate artifacts from padding regions at the end of the audio samples; AudioMAE effectively discards these regions, whereas AST does not. Except for SPC-2, which consists of a single word, there are two patch groups on either side of the diagonal in each histogram. The retention of these groups after pruning indicates the audio model’s reliance on (1) low-to-mid intensity patches with greater complexity and (2) high-intensity patches with less complexity. For general audio classification tasks (AS-20K, ESC-50), we observe that AudioMAE retains significantly more low-intensity tokens compared to AST. AudioMAE retains approximately 1.9× more tokens belonging to low-intensity clusters (C1 and C2) after pruning than AST in both datasets. While both models and datasets retain empty (bottom-left most) patches after pruning, AudioMAE prunes all of those patches in the speaker identification task (VoxCeleb-1).
IV Conclusion

### v2

@@ 4.5 Inspection of Retained Patches’ Statistics
5 Ablation Studies
5.1 Impact of Removing Tokens from Low or High Intensity Regions
We assess the impact of discarding tokens belonging to low- (C1, C2: ) or high-intensity (C4, C5: ) groups on accuracy during inference. In this test, the tokens belonging to these groups are pruned after being processed by specific blocks. In Table 6, and indicate the block index at which low- and high-intensity token groups are removed, respectively. High-intensity tokens contribute significantly to model performance. In addition, low-intensity tokens become more important for general audio classification tasks than for speech tasks. By comparing the gaps in accuracy drops between low-intensity and high-intensity tokens, we see that high-intensity tokens require deeper blocks to extract features for prediction. This also indicates that they contain richer information than low-intensity tokens. Overall, we could also observe that AudioMAE shows greater robustness to the loss of both token types across all datasets compared to AST.
5.2 Using Different Feature Aggregation Methods
We investigated whether different attention scoring and aggregation strategies—token-to-token attention with mean pooling versus [CLS] attention—affect the retention of low-intensity tokens, given that AudioMAE retains more such tokens than AST. We trained AudioMAE with a [CLS] token (AudioMAE-CLS) and AST with mean pooling (AST-MP) for classification. Fig. 8 compares the cumulative distribution of retained tokens with respect to mean patch intensity. AudioMAE retained more low-intensity tokens than AST and achieved higher accuracy in SPC-2 task. This difference can be explained by the models’ distinctive pretraining objectives: AudioMAE is trained in a self-supervised fashion to reconstruct every audio patch, regardless of its intensity. However, AST is trained in a supervised fashion, requiring the model to attend to tokens that are directly related to the label. Although AudioMAE does not outperform AST on the ESC-50 task, it appears to learn better audio representations, as evidenced by its superior accuracy on VoxCeleb-1.
6 Conclusion
@@
7 Discussions
Other pruning methods such as DynamicViT, which uses MLP modules as token score predictor, may exhibit different pruning patterns. We believe that the effectiveness of token pruning varies across tasks due to two factors, which we leave for future investigation: (1) how uniquely an audio model produces token representations, and (2) the number of tokens a task requires to capture audio features.